a pile of sand is in the shape of a cone. if the pile is 4 meters high and has a diameter of 7 meters, how many cubic meters of sand is in a pile?
Answer by checkley75(3666)

V=1/3BH
V=1/3(PI3.5^2)*4 (PI=3.14)
V=1/3*3.14*12.25*4
V=(3.14*12.25*4)/3
V=153.86/3
V=51.29 CUBIC METERS OF SAND.